# GAN Based Near-Field Channel Estimation for Extremely Large-Scale MIMO Systems

### v1

@@ III-D IE-Pix2pix Based Channel Estimation
2) Online prediction phase: In the online prediction phase, the trained generator in the IE-Pix2pix model is applied to the near-field channel estimation problem in XL-MIMO systems. After the offline training, the new initially estimated near-field channels are input into the trained generator in the IE-Pix2pix to directly produce XL-MIMO near-field channels composed of both LoS and NLoS path components.
@@ III-E Complexity Analysis
The computational complexity of the proposed IE-Pix2pix comes from two parts, i.e., initial estimation in equation (10) and the designed Pix2pix. The complexity of the initial estimation in equation (10) can be represented by $N_{\mathrm{t}}N_{\mathrm{r}}\left(N_{\mathrm{t}}+N_{\mathrm{r}}\right)$. The complexity of the designed Pix2pix comes from two CNNs, i.e., the generator and the discriminator. According to [35], the complexity of the Pix2pix can be expressed as $\mathcal{O}\left(\sum\nolimits_{s=1}^{S}{F_{s,h}F_{s,w}K_{s}^{2}N_{s}N_{s-1}}\right)$, where $K_{s}$ and $N_{s}$ represent the filter size and the number of output feature map in the $s$-th layer, respectively, $F_{s,w}$ and $F_{s,h}$ denote the width and the height of the feature map in the $s$-th layer, respectively. According to the aforementioned offline training of the IE-Pix2pix, the related parameters are listed in Table II, when $N_{\mathrm{t}}=256$ and $N_{\mathrm{r}}=128$. Therefore, the computational complexity of the proposed IE-Pix2pix can be expressed as
@@
IV Numerical Results
IV-A Simulation Setting
We consider the XL-MIMO near-field system, where the transmitter is equipped with $N_{\mathrm{t}}=256$ antennas and the receiver is equipped with $N_{\mathrm{r}}=128$ antennas. In this paper, we adopt the mixed LoS/NLoS XL-MIMO near-field channel model proposed in [22] and mainly focus on the channel estimation in the mixed LOS/NLOS near-field region. In this case, the distance between the transmitter and the receiver is smaller than the ARD, which can be calculated as $\mathrm{ARD}=\frac{4D_{\mathrm{t}}D_{\mathrm{r}}}{\lambda}\approx 197$ m. The carrier frequency is set as $f_{\mathrm{c}}=50$ GHz, which corresponds to $\lambda=0.006$ m. The XL-MIMO near-field channel in equation (3) is composed of $L=3$ NLoS path components. The angle $\theta_{\mathrm{t},l}$, angle $\theta_{\mathrm{r},l}$, complex gain $\alpha_{l}$, distance $d_{l}^{\mathrm{t}}$, and distance $d_{l}^{\mathrm{r}}$ are generated as following: $\theta_{\mathrm{t},l}\sim\mathcal{U}\left(-\frac{\pi}{3},\frac{\pi}{3}\right)$, $\theta_{\mathrm{r},l}\sim\mathcal{U}\left(-\frac{\pi}{3},\frac{\pi}{3}\right)$, $\alpha_{l}\sim\mathcal{C}\mathcal{N}\left(0,1\right)$, $d_{l}^{\mathrm{r}}\sim\mathcal{U}\left(10,130\right)$, and $d_{l}^{\mathrm{r}}\sim\mathcal{U}\left(10,130\right)$. Each element in the combining matrix $\mathbf{W}$ and the pilot matrix $\mathbf{Q}$ is randomly selected from $\left\{-\frac{1}{\sqrt{N_{\mathrm{r}}}},+\frac{1}{\sqrt{N_{\mathrm{r}}}}\right\}$ and $\left\{-\frac{1}{\sqrt{P}},+\frac{1}{\sqrt{P}}\right\}$, respectively.
According to the XL-MIMO near-field channel model in equation (9), we generate channel samples to train and test the proposed IE-Pix2pix. Once the channel samples are generated, the corresponding received signals and initially estimated channel samples can be obtained by using equations (2) and (10), respectively. The training set and test set consist of 4000 and 1000 samples, respectively. The optimizer for the generator and the discriminator is RMSProp. The learning rates for these two CNNs are 0.00002 and 0.0002, respectively. The batch size is 1.
@@ IV-B IE-Pix2pix Based Channel Estimation
Figure 7: NMSE versus SNR for different methods.
@@
We also test the NMSE performance of the proposed IE-Pix2pix under SNR $=10$ dB when the IE-Pix2pix is used to estimate the XL-MIMO near-field channel only including the LoS path component or NLoS path components. Fig. 6 shows the training process of the IE-Pix2pix to only estimate the LoS path component or NLoS path components. When the XL-MIMO near-field channel only includes the LoS path component, the proposed IE-Pix2pix has very good NMSE performance. Furthermore, the NMSE performance of the IE-Pix2pix to only estimate the LoS path component degrades with the increase in the distance $r$. This is because the LoS path component of the XL-MIMO near-field channel has different features with the change of the distance $r$. The IE-Pix2pix has very bad NMSE performance when the IE-Pix2pix is used to estimate the XL-MIMO channel only containing the NLoS path components. Hence, the bad performance of the IE-Pix2pix is mainly due to the large estimation error of the NLoS path components. Note that the objective of the proposed IE-Pix2pix is to simultaneously estimate the LoS and NLoS path components. Although the IE-Pix2pix has very bad performance when it is used to estimate the XL-MIMO channel only including the NLoS path components, the IE-Pix2pix can still achieve satisfactory NMSE performance in the adopted channel model when the distance $r$ is small. This is because the IE-Pix2pix can still efficiently learn the mapping from the initially estimated channels to the XL-MIMO near-field channels composed of the LoS and NLoS path components.
@@
Fig. 7 shows the impact of the distance $r$ on the NMSE performance when SNR $=$ 10 dB. The size of the pilot matrix is $256\times 128$. As the distance $r$ decreases, the NMSE performance of the proposed IE-Pix2pix improves gradually while the NMSE performance of the far-field codebook based OMP scheme degrades. This is because there are phase discrepancies between the far-field array response vectors based channel and the adopted channel [22], which can be learned by the IE-Pix2pix efficiently. The proposed IE-Pix2pix achieves better NMSE performance than the two stage algorithm when $r\leqslant$ 50 m. Meanwhile, the IE-Pix2pix performs better than the far-field codebook based OMP method when $r\leqslant$ 85 m. Furthermore, the IE-Pix2pix can even outperform the CRLB when the distance $r$ is enough small. In addition, both the CRLB and the two stage algorithm achieve almost unchanged NMSE performance when the distance $r$ varies. Note that the CRLB is the best unbiased estimator but not the greatest lower bound.
@@ V Conclusion
In this paper, we proposed a channel estimation method based on the IE-Pix2pix for XL-MIMO communication systems. In particular, the Pix2pix was developed to simultaneously estimate the LoS and NLoS path components of the XL-MIMO near-field channel. Moreover, we combined the adversarial loss introduced by the GAN architecture with a traditional loss function to ensure the right direction of training the generator. We also investigated the influence of the hyper-parameter of the loss function on the NMSE performance of the proposed method. To estimate the XL-MIMO near-field channels more accurately, the initially estimated channels instead of the received signals were fed into the designed Pix2pix. Numerical results showed that the proposed IE-Pix2pix based channel estimation method outperformed the two stage algorithm and the far-field codebook based OMP scheme in the adopted mixed LoS and NLoS XL-MIMO near-field channel model. Furthermore, the IE-Pix2pix surpassed the CRLB when the distance between the transmitter and the receiver was small and the pilot overhead was low.

### v2

@@ III-D IE-Pix2pix Based Channel Estimation
2) Online prediction phase: In the online prediction phase, the trained generator in the IE-Pix2pix model is applied to the near-field channel estimation problem in XL-MIMO systems. After the offline training phase, we can obtain the trained generator in the IE-Pix2pix. Then, the trained generator can be used to perform near-field channel estimation by using the new initially estimated near-field channels $\mathbf{H}_{\mathrm{IE}}$ in the online prediction phase. The new initially estimated near-field channels $\mathbf{H}_{\mathrm{IE}}$ are input into the trained generator in the IE-Pix2pix to output the estimated near-field channels $\hat{\mathbf{H}}$ composed of both the LoS and NLoS path components.
@@ III-E Complexity Analysis
The computational complexity of the proposed IE-Pix2pix comes from two parts, i.e., initial estimation in equation (10) and the designed Pix2pix. The complexity of the initial estimation in equation (10) can be represented by $N_{\mathrm{t}}N_{\mathrm{r}}\left(N_{\mathrm{t}}+N_{\mathrm{r}}\right)$. The complexity of the designed Pix2pix comes from two CNNs, i.e., the generator and the discriminator. According to [36], the complexity of the Pix2pix can be expressed as $\mathcal{O}\left(\sum\nolimits_{s=1}^{S}{F_{s,h}F_{s,w}K_{s}^{2}N_{s}N_{s-1}}\right)$, where $K_{s}$ and $N_{s}$ represent the filter size and the number of output feature map in the $s$-th layer, respectively. According to the aforementioned offline training of the IE-Pix2pix, the related parameters are listed in Table II, when $N_{\mathrm{t}}=256$ and $N_{\mathrm{r}}=128$. Therefore, the computational complexity of the proposed IE-Pix2pix can be expressed as
@@
IV Numerical Results
IV-A Simulation Setting
We consider the XL-MIMO near-field system, where the transmitter is equipped with $N_{\mathrm{t}}=256$ antennas and the receiver is equipped with $N_{\mathrm{r}}=128$ antennas. In this paper, we adopt the mixed LoS/NLoS XL-MIMO near-field channel model proposed in [24] and mainly focus on the channel estimation in the mixed LoS/NLoS near-field region. In this case, the distance between the transmitter and the receiver is smaller than the ARD, which can be calculated as $\mathrm{ARD}=\frac{4D_{\mathrm{t}}D_{\mathrm{r}}}{\lambda}\approx 197$ m. The carrier frequency is set as $f_{\mathrm{c}}=50$ GHz, which corresponds to $\lambda=0.006$ m. The XL-MIMO near-field channel in equation (3) is composed of $L=3$ NLoS path components. The angle $\theta_{\mathrm{t},l}$, angle $\theta_{\mathrm{r},l}$, and complex gain $\alpha_{l}$ are generated as following: $\theta_{\mathrm{t},l}\sim\mathcal{U}\left(-\frac{\pi}{3},\frac{\pi}{3}\right)$, $\theta_{\mathrm{r},l}\sim\mathcal{U}\left(-\frac{\pi}{3},\frac{\pi}{3}\right)$, and $\alpha_{l}\sim\mathcal{C}\mathcal{N}\left(0,1\right)$. Similar to [24], both the distance $d_{l}^{\mathrm{t}}$ and distance $d_{l}^{\mathrm{r}}$ are generated in range of $[10,130]$ meters. Each element in the combining matrix $\mathbf{W}$ and the pilot matrix $\mathbf{Q}$ is randomly selected from $\left\{-\frac{1}{\sqrt{N_{\mathrm{r}}}},+\frac{1}{\sqrt{N_{\mathrm{r}}}}\right\}$ and $\left\{-\frac{1}{\sqrt{P}},+\frac{1}{\sqrt{P}}\right\}$, respectively.
According to the XL-MIMO near-field channel model in equation (9), we generate channel samples to train and test the proposed IE-Pix2pix. Once the channel samples are generated, the corresponding received signals and initially estimated channel samples can be obtained by using equations (2) and (10), respectively. The training set and test set consist of 4000 and 1000 samples, respectively. The proposed IE-Pix2pix is implemented in Python 3.7.16 with Tensorflow 2.2.0. The optimizer for the generator and the discriminator is RMSProp. The learning rates for these two CNNs are 0.00002 and 0.0002, respectively. The batch size is 1.
@@ IV-B IE-Pix2pix Based Channel Estimation
Figure 7: NMSE performance against the distance for different methods.
@@
We also test the NMSE performance of the proposed IE-Pix2pix under SNR $=10$ dB when the IE-Pix2pix is used to estimate the XL-MIMO near-field channel only including the LoS path component or NLoS path components. Fig. 6 shows the training process of the IE-Pix2pix to only estimate the LoS path component or NLoS path components. When the XL-MIMO near-field channel only includes the NLoS path components, the proposed IE-Pix2pix achieves very bad NMSE performance. However, the IE-Pix2pix has very good NMSE performance when the IE-Pix2pix is used to estimate the XL-MIMO near-field channel only containing the LoS path component. Furthermore, the NMSE performance of the IE-Pix2pix to only estimate the LoS path component improves with the decrease in the distance $r$. This is because the LoS path component of the XL-MIMO near-field channel has different features with the change of the distance $r$ [24]. Hence, the NMSE performance of the IE-Pix2pix to estimate the XL-MIMO near-field channel including both the LoS and NLoS path components degrades gradually with the increase in the distance $r$, as shown in Fig. 5 and Fig. 7. The bad performance of the IE-Pix2pix to simultaneously estimate the LoS and NLoS path components is mainly due to the large estimation error of the NLoS path components. Note that the objective of the proposed IE-Pix2pix is to simultaneously estimate the LoS and NLoS path components. Although the IE-Pix2pix has very bad performance when the IE-Pix2pix is used to estimate the XL-MIMO channel only including the NLoS path components, the IE-Pix2pix can still achieve satisfactory NMSE performance in the adopted mixed LoS/NLoS channel model when the distance $r$ is small. This is because the IE-Pix2pix can still efficiently learn the mapping from the initially estimated channels to the XL-MIMO near-field channels composed of the LoS and NLoS path components.
@@
Fig. 7 shows the impact of the distance $r$ on the NMSE performance when SNR $=$ 10 dB. The size of the pilot matrix is $256\times 128$. As the distance $r$ decreases, the NMSE performance of the proposed IE-Pix2pix improves gradually while the NMSE performance of the far-field codebook based OMP scheme degrades. This is because there are phase discrepancies between the far-field array response vectors based channel and the adopted channel [24], which can be learned by the IE-Pix2pix efficiently. The proposed IE-Pix2pix achieves better NMSE performance than the two stage algorithm when $r\leqslant$ 50 m. The IE-Pix2pix without the IE can still outperform the far-field codebook based OMP method when $r\leqslant$ 80 m, while the IE-Pix2pix performs much better than the IE-Pix2pix without the IE. This result demonstrates that the initially estimated channels can help the proposed IE-Pix2pix generate the XL-MIMO near-field channels more efficiently. For the IE-Pix2pix without the IE, the strides in the first two layers of the generator need to be changed to rescale the received signals $\mathbf{Y}$ into the same size as the XL-MIMO near-field channels $\mathbf{H}$, since $\mathbf{Y}$ and $\mathbf{H}$ are of different sizes. Furthermore, the IE-Pix2pix can even outperform the CRLB when the distance $r$ is enough small. In addition, both the CRLB and the two stage algorithm achieve almost unchanged NMSE performance when the distance $r$ varies. Note that the CRLB is the best unbiased estimator but not the greatest lower bound.
@@ V Conclusion
In this paper, we proposed a channel estimation method based on the IE-Pix2pix for XL-MIMO communication systems. In particular, the Pix2pix was developed to simultaneously estimate the LoS and NLoS path components of the XL-MIMO near-field channel. Moreover, we combined the adversarial loss introduced by the GAN architecture with a traditional loss function to ensure the right direction of training the generator. We also investigated the influence of the hyper-parameter of the loss function on the NMSE performance of the proposed method. To estimate the XL-MIMO near-field channels more accurately, the initially estimated channels instead of the received signals were fed into the designed Pix2pix. Numerical results showed that the proposed IE-Pix2pix based channel estimation method outperformed the two stage algorithm and the far-field codebook based OMP scheme in the adopted mixed LoS/NLoS XL-MIMO near-field channel model. Furthermore, the IE-Pix2pix surpassed the CRLB when the distance between the transmitter and the receiver was small and the pilot overhead was low.